# Taming Data Challenges in ML-based Security Tasks: Lessons from Integrating Generative AI

### v1

@@ IX Conclusion & Future Work
We now address Section III-B questions using Section VII findings and suggest future directions.
@@
Challenges with instantiating GenAI models.    Unlike traditional data augmentation schemes (e.g., SMOTE), GenAI methods are highly complex ML schemes, which are challenging to correctly instantiate for wide diversity of security tasks. 5/6 GenAI schemes fail to generate data for at least 1/7 security tasks. TabDDPM suffered from mode collapse, TVAE failed to generate sufficient samples while class-conditioning, and TabSyn is unable to manage large, high-dimensional datasets. While REaLTabFormer shows potential in Tor task, both GReaT and REaLTabFormer fail on most tasks (6/7 and 5/7, respectively) due to limitations with high-dimensional feature vectors. Re-engineering LLM-based approaches like GReaT and REaLTabFormer for greater scalability with large, high-dimensional security datasets offers several benefits: (1) LLMs have robust methods for controlled text generation which could help to design effective sample-conditioning schemes, (2) potential for prompt-based data generation via instruction-tuning [24], (3) ability to handle complex data types (discrete or continuous) as everything is represented in text, and (4) ability to augment in overlapping distributions.
@@
GenAI-based data augmentation to boost classifier performance.   Data augmentation using GenAI is a promising approach to boost performance of diverse security classifiers—we improved performance for 5/7 security tasks, without any changes to the underlying classification algorithm. Therefore, a carefully engineered GenAI tool can reignite progress for security problems stagnated by slow algorithmic innovations—i.e., by simply addressing any underlying data challenges. Future work can explore more challenging scenarios, including overlapping class distributions, noisy labels, sparse feature vectors. Developments in LLMs [103, 64], Diffusion models [82, 74, 87], VAE [94, 34, 56] and GANs [54] such as strategic sampling methods and weighted learning approaches can be leveraged to produce rare/limited data points from training data manifolds.
@@
Impact of highly controlled generation.    Existing GenAI schemes do not provide mechanisms for highly controlled generation. To fill this gap, we proposed Nimai-S, a novel scheme based on a VAE using a discrete latent space to enable generation of samples targeting specific regions of the data distribution. This sample-conditioned approach achieved the best performance in 3 security tasks with complex biases. Future work can explore this direction using other schemes, such as LLMs, Diffusion models and GANs.
@@
Rapid recovery from concept drift.   We showed that GenAI can help maintain classifier performance post-deployment through a malware classification task exhibiting concept drift. A hybrid Nimai scheme, combining sample- and class-conditioning, enabled rapid recovery with a few labeled samples from the drifted set. While effective, our recovery approach has room for improvement. Future work could explore using GenAI to recover from adversarial attacks. We discuss the ethical considerations of this work in Appendix Section -A.

### v2

@@ 9. Conclusion & Future Work
We now address questions in Section 3.2 using our findings in Section 7 and also suggest future directions.
@@
Challenges with instantiating GenAI models.    Unlike traditional data augmentation schemes (e.g., SMOTE), GenAI methods are highly complex ML schemes, which are challenging to correctly instantiate for wide diversity of security tasks. 5/6 GenAI schemes fail to generate data for at least 1/7 security tasks. TabDDPM suffered from mode collapse, TVAE failed to generate sufficient samples while class-conditioning, and TabSyn is unable to manage large, high-dimensional datasets. While REaLTabFormer shows potential in Tor task, both GReaT and REaLTabFormer fail on most tasks (6/7 and 5/7, respectively) due to limitations with high-dimensional feature vectors. Re-engineering LLM-based approaches like GReaT and REaLTabFormer for greater scalability with large, high-dimensional security datasets offers several benefits: (1) LLMs have robust methods for controlled text generation which could help to design effective sample-conditioning schemes, (2) potential for prompt-based data generation via instruction tuning (Chung2022ScalingIL), (3) ability to handle complex data types (discrete or continuous) as everything is represented in text, and (4) ability to augment in overlapping distributions.
@@
GenAI-based data augmentation to boost performance.   Data augmentation using GenAI is a promising approach to boost performance of diverse security classifiers—we improved performance for 5/7 security tasks, without any changes to the underlying classification algorithm. Therefore, a carefully engineered GenAI tool can reignite progress for security problems stagnated by slow algorithmic innovations—i.e., by simply addressing any underlying data challenges. Future work can explore more challenging scenarios, including overlapping class distributions, noisy labels, sparse feature vectors. Developments in LLMs (wang2024llmautoda; li-etal-2024-search), Diffusion models (samuel2024generating; park2025raretofrequent; shao2024diffult), VAE (stocksieker2024data; fajardo2018vos; kim2024tvariational) and GANs (khorram2024taming) such as strategic sampling methods and weighted learning approaches can be leveraged to produce rare/limited data points from training data manifolds.
@@
Impact of highly controlled generation.    Existing GenAI schemes do not provide mechanisms for highly controlled generation. To fill this gap, we proposed Nimai-S, a novel scheme based on a VAE using a discrete latent space to enable generation of samples targeting specific regions of the data distribution. This sample-conditioned approach achieved the best performance in 3 security tasks with complex biases. Future work can explore approaches like LLMs, Diffusion models, and GANs.
@@
Rapid recovery from concept drift.   We showed that GenAI can help maintain classifier performance post-deployment through a malware classification task exhibiting concept drift. A hybrid Nimai scheme, combining sample- and class-conditioning, enabled rapid recovery with a few labeled samples from the drifted set. While effective, our recovery approach has room for improvement. Future work could explore using GenAI to recover from adversarial attacks.
We discuss the ethical considerations of this work in Section A (Appendix).
Acknowledgements.
This material is based upon work supported by the National Science Foundation under grant award numbers 2453818, 2453819 and 2453820. Neal Mangaokar was supported by an NSF Graduate Fellowship. The views and findings solely reflect those of the authors and should not be seen as representative opinions of any funding agency, given the absence of any financial conflicts.
Appendix A Ethical Considerations
We used data and models that have been publicly shared for research purposes. We did not use human subjects in our research. Team members were not exposed to sensitive, private information, or disturbing content. All experiments were conducted in controlled lab settings and no deployed services were affected by our newly trained classifiers. For the study investigating data challenges, we collected titles and abstracts of papers published between 2017 and 2021 from publicly available proceedings on the USENIX and ACM websites, ensuring full compliance with their respective terms and conditions. The proposed/studied GenAI schemes generate synthetic data in the feature space (used by the defense schemes) and not in the problem space. We do not propose any methods for translating the data from the feature space to the problem space, and doing so would be challenging. Therefore, this limits any potential dual-use concerns of such GenAI technology by adversaries. However, an adversary can use GenAI schemes to generate training data to build a surrogate classifier, which can potentially be used to craft adversarial samples against the defense classifier. However, we believe that the benefits of the work towards cybersecurity defenses outweigh any potential harm caused by our study.
Appendix B Paper selection methodology
In a semi-automated process, we shortlisted 35 cybersecurity papers (from top conferences) that use ML-based classifiers for network and application security tasks. We scraped titles and abstracts of 1,254 papers from USENIX Security and IEEE S&P (2017–2021), then ranked them using latent semantic indexing (deerwester1990indexing) based on manually curated keywords (e.g., classifier, prediction, accuracy, epoch). We manually reviewed top-ranked papers, selecting 25 that used security classifiers. An additional 10 papers came from team member suggestions.
Domain
Papers with task descriptions
Network
security
Network intrusion detection system (sikder20176thsense; mirsky5kitsune), infections with domain generation algorithm (schuppen2018fanci), botnets (jan2020throwing; zhu2021botspot++), IoT cryptojacking  (tekiner2022lightweight), telephony scams (peeters2018sonar; li2018machine), BGP serial hijackers (testart2019profiling), malicious traffic (holland2021new; yu2020you), abnormal cyber-physical patterns (chen2018learning), and covert tunneling (barradas2018effective).
Social
media
& web
security
Impersonation (cidon2019high), phishing emails (ho2019detecting), abusive social media accounts (xu2021deep), stock pump-and-dump schemes (xu2019anatomy), web traffic/ad fraud (yang2022wtagraph), malicious URLs (thirumuruganathan2022siraj; chen2021cost), website fingerprinting (gong2022surakav), GDPR violations in cookies (bollinger2022automating), and security vs. non-security browser crash causes (park2019empirical).
Software
security
Malicious files (chen2020training), executable binary code (jordaney2017transcend; downing2021deepreflect; yang2021cade; barbero2022transcending; yang2021bodmas), software crash analysis (guo2019deepvsa), resilience against automated attacks (banescu2017predicting; han2021sigl; she2020neutaint), and vulnerability detection (she2019neuzz; zhou2019devign).
Table 6. Measurement study papers with task categories.
  $\Delta G$ for months
Technique  1 4 8 12
  Nimai-S  -8.3 (3.9) $\downarrow$ -5.4 (7.1) - -12.1 (10.1) $\downarrow$ -7.6 (6.7) $\downarrow$
Nimai-C  -8.8 (3.6) $\downarrow$ 6.5 (7.4) - 6.0 (9.2) - -12.5 (9.1) $\downarrow$
TVAE  -12.5 (5.3) $\downarrow$ -3.3 (10.1) - -4.8 (18.6) - -35.6 (10.2) $\downarrow$
TabSyn  -11.4 (4.7) $\downarrow$ -16.8 (9.5) $\downarrow$ -25.2 (6.0) $\downarrow$ -34.6 (14.6) $\downarrow$
CTAB-GAN+  -4.9 (4.4) $\downarrow$ 3.5 (10.0) - 3.6 (14.9) - -8.3 (15.2) -
TabDDPM  -2.5 (0.8) $\downarrow$ -2.5 (0.8) $\downarrow$ -1.4 (4.4) - -12.1 (7.1) $\downarrow$
SMOTE  -2.9 (1.1) $\downarrow$ -2.9 (2.1) $\downarrow$ -6.0 (5.2) $\downarrow$ -9.6 (7.4) $\downarrow$
MC-CCR  -1.5 (0.6) $\downarrow$ -3.4 (1.6) $\downarrow$ -5.7 (5.9) - -19.5 (14.4) $\downarrow$
Table 7. Mean $\Delta G$ with standard deviation for BODMAS task for 1, 4, 8, & 12 months. $\downarrow$ indicates negative gains. ‘-’ indicates unreliable gains (CV $\geq$ 1). None of the methods show positive $\Delta G$ in these months. GReaT and REaLTabFormer’s failures are not shown.
Task  TVAE CT TD TabSyn GReaT RT Nimai
  BGP  1 8.2 0.004 18.6 5.9 14 0.6
Tor  2 168 0.158 209 0.92 d* 1041 9.46
IoT  16.6 837 0.27 652 238 d* 2.7 h* 72
Cookie  77.5 2,074 0.223 51 $\times$ 1.45 h* 16.5
MS  10 206 0.01 $\times$ $\times$ $\times$ 3.57
BODMAS  3.9 232 0.123 14.7 $\times$ $\times$ 1
nPrintML  59.4 2,368 0.047 $\times$ $\times$ $\times$ 21.8
  Avg. time  24.3 841.8 0.119 - - - 17.84
Table 8. Training times per epoch for GenAI schemes on an NVIDIA A100 GPU (in seconds unless marked as h = hours or d = days). $\times$ indicates model failure; * indicates estimated time. CT, TD, and RT refer to CTAB-GAN+, TabDDPM, and REaLTabFormer, respectively.
Appendix C Implementation of GenAI/non-GenAI tools
This section discusses the hyper-parameters and implementation details of our chosen GenAI/non-GenAI tools.
TVAE (tvae/ctgan).    We use Synthetic Data Vault (SDV) (sdv) to implement TVAE with early stopping. The model has 7 hyper-parameters: latent space size, encoder/decoder units, learning rate, batch size, loss factor, and number of epochs. As the authors provide no tuning guidelines, we use default settings.
CTAB-GAN+ (ctabgan+).   We implement early stopping on CTAB-GAN+’s official source code. CTAB-GAN+ has 7 hyper-parameters, which include number of classifier layers, size of classifier layers, batch size, number of training epochs, number of channels, learning rate, number of encoding sides. Similar to TVAE, the authors provide no guidelines on how to tune the hyper-parameters, so we use the default settings.
TabDDPM (tabddpm).   We implement TabDDPM using the official codebase. It has 10 hyper-parameters, including learning rate, iterations, diffusion timesteps, batch size, and MLP layers. The authors recommend and use the Tree-Structured Parzen Estimator (watanabe2023tree) (TPE) for hyper-parameter search.
TabSyn (tabsyn).   We implement TabSyn using the official code, adding early stopping to the VAE training (already built-in for the Diffusion model). TabSyn’s architecture has 7 hyper-parameters: for the VAE—token dimension, number of layers, transformer layer dimension, $\lambda$, and min/max $\beta$; for the Diffusion model—hidden dimension. We skip hyper-parameter search, as the authors claim their parameter settings generalize across feature sizes.
GReaT (GreaT).   We use GReaT’s Python package with distilGPT as the backend LLM. We add early stopping and increase the context window from 512 to 1024 for tasks with $\geq$100 features. GReaT has no tunable hyper-parameters.
REaLTabFormer (solatorio2023realtabformer).   We implemented REaLTabFormer using their official Python package.
SMOTE (smote).    We use the imbalanced-learn (imbalanced-learn) implementation of SMOTE, with the number of neighbors in the $k$-nearest neighbors algorithm set to the default of 5.
MC-CCR (mc-ccr).   We implemented MC-CCR using their source code, and best recommended generation settings: proportional selection and translation cleaning strategy.
Appendix D DeepSMOTE methodology
DeepSMOTE (dablain2022deepsmote) uses an encoder-decoder architecture and SMOTE like interpolation in the latent space. During training, feature space samples are encoded into the latent space, where they are permuted and decoded to learn varied sample generation. During generation, feature space samples are encoded to the latent space, where SMOTE is applied, and decoded to produce feature space synthetic samples.
Task BGP Tor IoT Cookie MS BODMAS nPrintML
# of
samples
146 5,749 41,634 12,882 619 11,592 686
Table 9. Number of synthetic samples added during augmentation to balance out class distribution for all tasks. Here, MS refers to MS-Malware.
Appendix E MC-CCR methodology
Here we describe MC-CCR’s methodology in detail. MC-CCR (mc-ccr) follows a two-step process. In the first step, a cleaning procedure is performed i.e., the majority samples in the proximity of the minority samples are removed. In the second step, additional minority samples are added in spherical regions around each minority sample. MC-CCR uses a weighting strategy similar to ADASYN (adasyn) to compute the positions of synthetic samples to be added around difficult samples. The augmentation radius (sphere) is determined by the energy or budget allocated.
Appendix F Nimai Hyper-parameter search methodology
Nimai’s hyper-parameter search is conducted in two stages: first for the VQVAE model, then for the MTM model. To speed up the process, we use a stratified 10k-sample subset for tasks with over 10k training samples. Table 10 lists the architectural hyper-parameters and their search spaces. For VQVAE, ASHA identifies parameters that minimize codebook (embedding + commitment) loss on a validation set, excluding those causing codebook collapse. For MTM, ASHA minimizes cross-entropy loss on the validation set.
Component
Hyper-parameter Search space
Encoder,
Decoder,
MTM
Number of multi-heads
in attention layer.
{2, 4, 8, 16}
Number of neurons in
feed-forward layer.
{16, 32, 64,
128, 256, 512}
Number of transformer
layers.
{1, 2, 4}
VQVAE
Number of vectors in
codebook.
{32, 64, 96, 128,
256, 512, 1024}
Length of vectors
in the latent space.
2^n where n
between (5, 49)
Size of embedding
vectors in codebook.
{2 , 4, 6, 12,
16, 24, 32, 48}
Weighting factor of
codebook loss.
(1, 50)
Weighting factor of
reconstruction loss.
(1, 50)
Decay loss factor in
VQVAE EMA quantizer.
(0, 1)
Table 10. Architectural hyper-parameters used in Nimai’s VQVAE, encoder/decoder and MTM model components.
(a) IoT
(b) BGP
Figure 5. Feature-wise skewness comparison of attack-class samples before and after augmentation.:
(N-1) Month
F-score for months
(w/ standard deviation)
Subset size
  5 6
  1  56.23 (4.55) / 35.4 52.26 (2.77) / 33.8 34
2  42.79 (2.81) / 36.4 37.79 (4.01) / 33.4 49
3  59.78 (8.34) / 34.4 53.92 (10.39) / 33.4 64
4  50.11 (5.10) / 39 44.13 (3.85) / 33.7 33
5  48.27 (5.20) / 39.6 9
Table 11. Mean macro F-scores w/ standard deviation for Nimai-hybrid separated (/) by macro F-scores for INSOMNIA on BODMAS months 5 & 6. Best scores are bolded.
Appendix G Monthwise statistical tests on BODMAS
This section provides a month-by-month analysis of the 8 testing months where positive gains are observed, presented in Table 3. In months 2, 3 and 11, Nimai-C is the only scheme that shows a positive gain over all other schemes (both GenAI and non-GenAI), further validating Finding 3 (Section 7.1). In months 5, 6, 7, 9, and 10, several schemes show positive gains with high std values. Therefore, we apply statistical tests to compare the rankings of reliable schemes for these months.
In months 5 and 6, all GenAI schemes show positive gains. These months also experience the highest performance degradation under concept drift (Figure 2). The Kruskal-Wallis test shows a significant difference in rankings in both months 5 and 6, with p-values of 1.044e-5 and 2.42e-3 (<< 0.05), respectively. Dunn’s test for month 5 ranks the schemes as follows: Nimai-C >> CTAB-GAN+ >> TVAE >> Nimai-S >> TabDDPM >> TabSyn; clearly showing Nimai-C to be the best performing scheme amongst all GenAI schemes with a percentage gain of 59.42%. Dunn’s test for month 6 ranks the schemes as follows: TVAE >> CTAB-GAN+ >> Nimai-S >> Nimai-C >> TabDDPM >> TabSyn, with TVAE outperforming all other schemes. Note, Nimai-S is the only sample-conditioned scheme to outperform TabDDPM, TabSyn and Nimai-C class-conditioned GenAI schemes. Traditional sample-conditioned schemes exhibit unreliable performance in both months.
In month 7, only Nimai-C, TabDDPM and SMOTE show positive gains. Kruskal-Wallis test shows a significant difference in means with a p-value of 0.0982 << 0.05. Post-hoc tests rank the schemes as SMOTE >> Nimai-C >> TabDDPM. Here, traditional sample- conditioned SMOTE method outperforms all class-conditioned schemes including Nimai-C and TabDDPM.
In month 9, Nimai-C and TVAE show positive gains. In month 10, Nimai-C, Nimai-S, and TabDDPM show positive gains. The results of Kruskal-Wallis test indicate that there is no difference in means for the schemes with positive gains in both months. p-values for months 9 and 10 are 0.409 and 0.239 (>>0.05) respectively. Therefore, in month 9, Nimai-C and TVAE are the top performers, and in month 10, Nimai-C, Nimai-S and TabDDPM lead, all 3 of them being class-conditioned schemes.
Appendix H Impact of noisy class labels
We study how noisy/overlapping class samples affect data augmentation. In the original BGP task, sample-conditioned schemes show positive gains (Table 2). We induce class overlap by flipping 20% of class-0 labels. Results for top-performing GenAI and traditional methods are in Table 12. All methods degrade under overlap, highlighting the limited robustness of current augmentation schemes, especially GenAI, to noisy/overlapping data.
Technique
Mean $\Delta G$
  BGP Overlap
  Real  40.50 (-)
Nimai-S  -14.02 (2.26) $\downarrow$
Nimai-C  -0.46 (2.28) $\downarrow$
CTAB-GAN+  -25.35 (2.54) $\downarrow$
SMOTE  -37.40 (3.99) $\downarrow$
Table 12. Mean $\Delta G$ with standard deviation for overlapping class data setting in BGP. $\downarrow$ indicates negative gains.
# Test samples for BODMAS months
1 2 3 4 5 6 7 8 9 10 11 12
826 829 1563 1205 623 1087 945 1073 283 268 312 559
Table 13. Test set statistics for BODMAS over 12 months.
Appendix I Nimai Training Objective
Nimai’s training objectives include the following loss terms.
Reconstruction loss ($\mathcal{L}_{\text{recon}}$): This term ensures accurate reconstruction in feature space using mean squared error (MSE). For input $x$ and reconstruction $\hat{x}$:
(3) $\mathcal{L}_{\text{recon}}=\|x-\hat{x}\|^{2}$
Commitment Loss ($\mathcal{L}_{\text{commit}}$): This aligns encoder outputs with the nearest codebook embeddings, where $z_{e}(x)$ is the encoder output, $e$ the closest embedding, $sg(\cdot)$ the stop-gradient operator, and $\beta$ a scaling term:
(4) $\mathcal{L}_{\text{commit}}=\beta\|z_{e}(x)-\text{sg}(e)\|^{2}$
Embedding Loss ($\mathcal{L}_{\text{embed}}$): This updates embedding vectors to match encoder outputs:
(5) $\mathcal{L}_{\text{embed}}=\|\text{sg}(z_{e}(x))-e\|^{2}$
Finally, the overall objective combines all losses, weighted by $\alpha$:
(6) $\mathcal{L}=\mathcal{L}_{\text{recon}}+\alpha\left(\mathcal{L}_{\text{embed}}+\mathcal{L}_{\text{commit}}\right)$
Appendix J MTM Loss Function and Training Objective
MTM’s objective in Equation 7 uses cross-entropy loss to estimate the log-likelihood of quantized latent tokens. Here, $\mathbf{y}_{i}$ is a latent feature, $\mathbf{Y}_{\text{mask}}$ a masked token filled by MTM, and $\mathbf{c}$ the class label.
(7) $\mathcal{L}_{\text{MTM}}=\mathbb{E}_{(\mathbf{y},c)\sim D}\left[-\sum_{i}\log p\left(\mathbf{y}_{i}\mid\mathbf{Y}_{\text{mask}},c\right)\right]$